OpenStudy (anonymous):

At a track meet, 50 people ran the 100-meter dash. 2 people finished in 11 seconds, 5 people finished in 12 seconds, 8 people finished in 13 seconds, 10 people finished in 14 seconds, 21 people finished in 15 seconds, 2 people finished in 16 seconds, and 2 people finished in 17 seconds. What is the probability distribution for the finish times?

OpenStudy (dumbcow):

do a histogram 11 sec: 2 people: probability is 2/50 = 1/25 = 0.04 12 sec: 5 people: probability is 5/50 = 1/10 = 0.1 13 sec: 8 people: probability is 8/50 = 4/25 = 0.16 ... and so on